OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (7, 1) m = 10 is it y=69?

OpenStudy (anonymous):

y = 10x - 69

OpenStudy (anonymous):

y-1=10(x-7) y=10x-69